Question 4




(a)


(i) Explain the theory of constraints and throughput accounting, including the five key steps advocated by Goldratt and Cox.


(5 marks)




(ii) Identify and describe the three key measures advocated by Goldratt and Cox.


(3 marks)




(iii) Define fully the throughput accounting ratio used by Galloway and Waldron.


(2 marks)




(b)


BHJ Ltd produces a single product X, which passes through three different processes, A, B and C. The throughput per hour of the three processes is 16, 8 and 16 units of X respectively. The company works an 8 hour day, 5 days a week, 45 weeks a year. The selling price of X is £140 per unit and its material cost is £35 per unit. Conversion costs are planned to be £23,000 per week.






You are required to:




(i) Determine the throughput accounting (TA) ratio per day. (4 marks)




(ii) Calculate how much the company could spend on equipment to improve the throughput of process B if it wished to recover its costs in the following time periods.




2 years




28 weeks (8 marks)






(iii) Calculate the revised TA ratio if this money is spent.


1
Expert's answer
2021-01-15T07:56:15-0500

Question 4

(a)

(i) Explain the theory of constraints and throughput accounting, including the five key steps advocated by Goldratt and Cox.

The theory of constraints provides approaches that could be utilized to identify a vital limiting factor that tend to jeopardize the process of achieving a specific business objective and goal and improving the constraint to reducing its limitation aspects in a project. On the same note, a throughput accounting is a strategy used in accounting and finance to emphasize in enhancing throughput without allocating costs, especially on investments and expenses. The five key steps advocated by Goldratt and Cox include, identifying the constraint, exploiting it, subordinating, elevating and preventing inertia by repeating the process.

(ii) Identify and describe the three key measures advocated by Goldratt and Cox.

The three key performance measures include throughput, inventory, operating expense. The throughput measurement emphasizes on the generation of money through sales but nit production. Inventory is the amount invested in commodities that an organization intends to distribute while operating expense entails all the money that a company spends.

(iii) Define fully the throughput accounting ratio used by Galloway and Waldron.

The throughput accounting ratio (TPAR) is the return per factory hour relative to the cost per factory hour. It could be summarized as: TPAR = Return per factory hour/cost per factory hour.

(b)

BHJ Ltd produces a single product X, which passes through three different processes, A, B and C. The throughput per hour of the three processes is 16, 8 and 16 units of X respectively. The company works an 8 hour day, 5 days a week, 45 weeks a year. The selling price of X is £140 per unit and its material cost is £35 per unit. Conversion costs are planned to be £23,000 per week.

You are required to:

(i)                Determine the throughput accounting (TA) ratio per day. (4 marks)

140 – 35 = 105

105/35 = 3

TA ratio = 3

(ii)              Calculate how much the company could spend on equipment to improve the throughput of process B if it wished to recover its costs in the following time periods.

2 years

B = 8 units *35 = 280

280*8hours*5days*45 weeks*2 years = £1,008,000

28 weeks

280*8*5*28 = £313,600